The supplement of an angle is 60 degrees less than twice the supplement of the complement of the angle. Find the measure of the complement.
Answer by BurningFlame_marc(77)

Hello!

First let's represent, 180-x=supplement of the angle, 90-x= complement of a angle.
Equation:
180-x=2[180-(90-x)]-60

Solution:
180-x=2(180-90+x)-60
x=360-180+2x-60-180
-x=-60
x=60

Final Answer: The complement is 30